# Daniel Iong, Matthew McAnear, Yuezhou Qu

### v1

@@ 4 Simulation Studies
This measure is commonly used to compare predictive performance on unseen data among different models [27]. For GPR-G and GPR-CN, $p_{\text{pred}}$ is given in eq. 14 and eq. 22, respectively. For GPR-t and GPR-L, $\log p_{\text{pred}}$ is approximated using Monte-Carlo methods, i.e.
@@
where $M$ is the number of Monte-Carlo samples; $q$ is the variational distribution given in eq. 13; and $\log p(y^{*}|f^{(m)},\vartheta)$ is the log-likelihood function for either the Student-t or Laplace distribution centered at $f^{(m)}$ with hyperparameters $\vartheta$. We set $M=1000$ when computing the NLPD for GPR-t and GPR-L. Furthermore, the expectation term in eq. 11 for GPR-t and GPR-L are also intractable and are approximated using the Gauss-Hermite quadrature method [28]. In each of the simulations below, we use the popular squared exponential kernel with automatic relevance determination (ARD) defined as
@@
where $\sigma_{s}^{2}$ is an output scale parameter; and $\{\ell_{j}\}_{j=1}^{p}$ are individual length scales for each input dimension. We train each SVGPR model with 500 inducing points for 30 iterations using the Adam algorithm implemented in PyTorch with an exponentially decaying step size starting at 0.1 and a batch size of 256 [29, 30]. The inducing inputs were initialized to be a random sample of the training inputs. Since the objective function for GPR-CN and GPR-t may have multiple local optima, we rerun our algorithm five times with different initial parameter values and keep the run that yields the largest value for eq. 18 in each simulation study.
@@ 4.1 Parameter estimation
$f(x_{i})=0.3+0.4x_{i}+0.5\text{sin}(2.7x_{i})+\frac{1.1}{1+x_{i}^{2}},$ (24)
where $x_{1},\dots,x_{5000}\sim\text{Uniform}(0,5)$. We plot one of the simulated datasets in fig. 2. This was adapted from an artificial regression problem considered by [31]. We plot the parameter estimates and corresponding ELBO value across iterations for one dataset in fig. 3. In this specific run, the parameter estimates converge to a value close to the true value after roughly 35 iterations. Figure 4 shows boxplots for estimated parameter values across the different simulated datasets. These boxplots show that most of the estimated values are close to the true parameter values. We plot the estimated mean functions and their respective RMSE values in fig. 5. From these plots, we can see that the estimated mean functions are close to the true function most of the time.
@@ 4.2 Comparison with other robust likelihoods
For the second simulation study, we consider an artificial regression problem described in [10] which uses the following function first introduced in [32]:
@@
Lastly, we add outliers by replacing $p_{\text{outlier}}$ proportion of the generated observations with samples drawn from $N(15,\sigma_{\text{outlier}}^{2})$, where $p_{\text{outlier}}\in\{0.1,0.2,0.3\}$ and $\sigma_{\text{outlier}}\in\{3,10\}$. In the $\sigma_{\text{outlier}}=3$ case, the generated outliers are likely to lie in the same range as the function values. Setting $\sigma_{\text{outlier}}=10$ constitutes a more difficult case where the outliers are unrelated to the function and are likely to lie outside of the function value range. We consider four scenarios with increasing outlier proportion and magnitude to study the performance of the considered noise models. These scenarios are summarized in table 1. The first scenario is the same one considered in [10] and serves as a baseline for the remaining scenarios which have an increasing proportion of extreme outliers. We fit the various GPR models to simulated datasets generated under various outlier scenarios and evaluate them on 10,000 noise-free test samples of eq. 25.
@@
Table 1: Summary of outlier scenarios considered in second simulation study.
Boxplots of the RMSE, MAE, and NLPD for the different models and outlier scenarios are given in figs. 6, 7, 8 and 9. Note that the boxplots for GPR-CN and GPR-t tend to have a higher spread compared to GPR-G and GPR-L. This is likely because these methods have ELBOs that are not log-concave and are vulnerable to local optima. Furthermore, the NLPDs for GPR-t are approximated with Monte-Carlo samples. In all scenarios, the robust noise models (GPR-CN, GPR-t, GPR-L) outperform GPR-G in terms of all metrics. The RMSEs and MAEs are similar for the robust noise models (GPR-CN, GPR-t, GPR-L) in all scenarios. In the first scenario, the NLPDs are similar for GPR-CN and GPR-t but are slightly higher for GPR-L. This difference in NLPDs is exacerbated in the remaining scenarios, where the outliers are more extreme. This is likely because the Laplace distribution is not as heavy-tailed as the t-distribution and does not model outliers separately as in the CN distribution. As we increase the proportion of outliers, GPR-CN starts to outperform GPR-t in terms of NLPD. This suggests that GPR-CN is more adept at handling larger proportions of outliers than GPR-t. This may be due to the fact that outliers are identified and placed into its own separate component in GPR-CN so they don’t affect model fit for non-outliers. On the other hand, outliers are not explicitly isolated in GPR-t so they may affect the fit for non-outliers. As the proportion of outliers increases, we may expect the true latent function posterior to be multimodal, making the Gaussian variational distribution inadequate for approximating it. However, this does not seem to hinder predictive performance as can be seen from the results in the third and fourth scenario.
Figure 6: $p_{\text{outlier}}=0.1$, $\sigma_{\text{outlier}}=3$. Boxplots of NLPD, RMSE, and MAE for second simulation study with varying $p_{\text{outlier}}$ and $\sigma_{\text{outlier}}$ across 200 simulated datasets.
Figure 7: $p_{\text{outlier}}=0.1$, $\sigma_{\text{outlier}}=10$. Boxplots of NLPD, RMSE, and MAE for second simulation study with varying $p_{\text{outlier}}$ and $\sigma_{\text{outlier}}$ across 200 simulated datasets.
Figure 8: $p_{\text{outlier}}=0.2$, $\sigma_{\text{outlier}}=10$. Boxplots of NLPD, RMSE, and MAE for second simulation study with varying $p_{\text{outlier}}$ and $\sigma_{\text{outlier}}$ across 200 simulated datasets.
Figure 9: $p_{\text{outlier}}=0.3$, $\sigma_{\text{outlier}}=10$. Boxplots of NLPD, RMSE, and MAE for second simulation study with varying $p_{\text{outlier}}$ and $\sigma_{\text{outlier}}$ across 200 simulated datasets.
@@ 5 Applications
In this section, we fit the four SVGPR models identified in section 4.2 on flight delays. For ground magnetic perturbations data ($\delta B_{H}$), we fit a dense artificial neural network and a GPR-CN model and compare their predictive performances with skill scores, RMSE, and focus in depth on invterval coverage.
The model fitting procedure varies slightly between datasets. For the flights data, we train the four GPR models with 1000 inducing points for up to 30 iterations using the Adam algorithm implemented in PyTorch with an exponentially step size starting at 0.1 and a batch size of 256 [29, 30]. The inducing inputs were initialized to be a random sample of the training inputs. We used a validation set to monitor the validation NLPD and terminate training if it hasn’t decreased in 5 iterations. Input features are standardized by subtracting the training sample mean and dividing by the training sample standard deviation.
For $\delta B_{H}$ prediction, we use a more flexible hyperparameter setup, opting to allow the number of inducing points and batch size to vary and choosing optimal hyperparameters using the Bayesian Optimization package [33]. As the data is fully numeric and relatively small station by station, we use a large batch size on a GPU to speed up computation. The details of this can be found in 5.2.2.
@@ 5.1 Flight delays
The flight delays dataset consists of information about commercial flights in the US from January 2008 to April 2008 and was taken from an example in [23]. It has become a standard benchmark dataset in the GPR literature for comparing scalable GPR methods due to its massive size and non-stationary nature. The original dataset curated by [23] contains 800,000 observations that were randomly sampled from a dataset containing around 2 million observations. To speed up computation, we reduced the size by randomly sampling 100,000 observations for training, and 20,000 each for validating and testing. The goal is to predict flight delay times (in minutes) using the eight features considered by [23]: aircraft age, distance to travel, airtime, departure time, arrival time, day of the week, day of the month, and month. A histogram of flight delay times used for training is given in fig. 10. 99% of flights are only delayed for up to 166 minutes. However, there are a handful of flights that are delayed for more than 300 minutes. These delays are likely due to external factors such as weather that are not considered in our models. This motivates the use of robust models for analyzing this data.
@@ 5.2.1 Data Setup
Table 4: All test data. Summary statistics of GPR-CN and ANN models for the testing period of 2015-2016. Target coverage is 95%. Coverage is comparable in both storm and non-storm periods, but the IQR of interval length is generally either comparable or lower in the GPR-CN model.
@@
Table 5: Storm periods only. Summary statistics of GPR-CN and ANN models for the testing period of 2015-2016. Target coverage is 95%. Coverage is comparable in both storm and non-storm periods, but the IQR of interval length is generally either comparable or lower in the GPR-CN model.
@@
Following the convention made in the space weather community, we do not include prior values of the output as inputs [34, 35, 36]. Therefore, we will ignore the temporal nature of this data and treat it as a conventional regression problem. Space physicists are typically most interested in the predictive performance during geomagnetic storms when $\delta B_{H}$ spikes, and so we report results across the entire prediction range as well as within storm performance.
@@ 5.2.2 Models
To fit the GPR-CN models, we use a Linux server with 32GB of RAM, a 16-core processor, and a 4GB NVidia GeForce 2060 GPU. As with the ANN, we use the Bayesian-Optimization library [33] to find optimal hyperparameters. While the GPU is not technically necessary, it should be noted that hyperparameter tuning on the CPU is a lengthy process, and GP models can be susceptible to the starting points of tuning parameters. The GPU enables running many more iterations of both model fitting and hyperparameter tuning, though it comes at a cost of model flexibility, namely by limiting the kernel setup. Due to memory constraints on the GPU, a periodic component could not be added to the existing kernel, and instead, seasonal features were added to the data to remedy this. Next, GPU training adds some complexity to the code primarily around data-type handling, though this is mitigated by the robust support of both the PyTorch and GPyTorch packages.
Though GPU-training is optional, it vastly improved the number of iterations we were able to perform for both hyperparameter training and model testing, and is therefore recommended for any extensions to this research.
@@ 5.2.3 Results
In terms of accuracy, the ANN model performs slightly better than the GPR-CN model. Overall RMSE is lower on most stations, though both models are comparable. In addition to RMSE, we calculate Heidke and True Skill scores using station-specific deciles, as in [37].
@@
The primary advantage of the GPR-CN model is the interval length. Though the conservatism of the GPR-CN model causes it to have marginally diminished accuracy compared to the ANN, it also constrains the model from making larger outlier predictions. The ANN often overestimates the standard deviation to the point of creating arbitrarily large confidence intervals that achieve coverage by guessing an impossibly large range. This problem is especially pronounced in storm periods.
Though the GPR-CN model sacrifices accuracy when compared to the ANN, it does so for the benefit of having better constrained prediction intervals. For nine stations of eleven, the median interval length was lower than the ANN, and in all stations, the GPR-CN model had lower maximum interval lengths for similar coverage (table 4, fig. 13). Further, as the RMSE of a particular station increases, the IQR of interval lengths increases as well, but this increase is less severe for the GPR-CN model (see fig. 11). This improvement to confidence intervals is especially pronouned examining the high RMSE (re: high latitude) stations during storm periods.
@@ 6 Conclusion
In this paper, we proposed a scalable inference algorithm for fitting sparse GPR models with contaminated normal noise. In section 4.1, we showed that the algorithm is able to correctly estimate the model hyperparameters when used to fit data simulated from the model itself. In section 4.2, we compared GPR models with different likelihoods and showed that our proposed model outperforms the others when there is a large proportion of outliers. In section 5, we applied our proposed algorithm to two real-world datasets and discussed strengths and weaknesses compared to GPR models with other robust likelihoods. In particular, our model was able to provide more informative prediction intervals for the ground magnetic perturbations dataset.

### v3

@@ 4 Simulation Studies
This measure is commonly used to compare predictive performance on unseen data among different models (Gelman et al., 2014). For GPR-G and GPR-CN, $p_{\text{pred}}$ is given in eq. 14 and eq. 22, respectively. For GPR-t and GPR-L, $\log p_{\text{pred}}$ is approximated using Monte-Carlo methods, i.e.
@@
where $M$ is the number of Monte-Carlo samples; $q$ is the variational distribution given in eq. 13; and $\log p(y^{*}|f^{(m)},\vartheta)$ is the log-likelihood function for either the Student-t or Laplace distribution centered at $f^{(m)}$ with hyperparameters $\vartheta$. We set $M=1000$ when computing the NLPD for GPR-t and GPR-L. Furthermore, the expectation term in eq. 11 for GPR-t and GPR-L are also intractable and are approximated using the Gauss-Hermite quadrature method (Liu and Pierce, 1994). In each of the simulations below, we use the popular squared exponential kernel with automatic relevance determination (ARD) defined as
@@
where $\sigma_{s}^{2}$ is an output scale parameter; and $\{\ell_{j}\}_{j=1}^{p}$ are individual length scales for each input dimension. We train each SVGPR model with 500 inducing points for 30 iterations using the Adam algorithm implemented in PyTorch with an exponentially decaying step size starting at 0.1 and a batch size of 256 (Kingma and Ba, 2014; Paszke et al., 2019). The inducing inputs were initialized to be a random sample of the training inputs. Since the objective function for GPR-CN and GPR-t may have multiple local optima, we rerun our algorithm five times with different initial parameter values and keep the run that yields the largest value for eq. 18 in each simulation study.
@@ 4.1 Parameter estimation
$Figures/f(x_{i})=0.3+0.4x_{i}+0.5\text{sin}(2.7x_{i})+\frac{1.1}{1+x_{i}^{2}},$ (24)
where $x_{1},\dots,x_{5000}\sim\text{Uniform}(0,5)$. We plot one of the simulated datasets in fig. 2. This was adapted from an artificial regression problem considered by Neal, (1997). We plot the parameter estimates and corresponding ELBO value across iterations for one dataset in fig. 3. In this specific run, the parameter estimates converge to a value close to the true value after roughly 35 iterations. Figure 4 shows boxplots for estimated parameter values across the different simulated datasets. These boxplots show that most of the estimated values are close to the true parameter values. We plot the estimated mean functions and their respective RMSE values in fig. 5. From these plots, we can see that the estimated mean functions are close to the true function most of the time.
@@ 4.2 Comparison with other robust likelihoods
For the second simulation study, we consider an artificial regression problem described in Kuss, (2006) which uses the following function first introduced in Friedman, (1991):
@@
Lastly, we add outliers by replacing $p_{\text{outlier}}$ proportion of the generated observations with samples drawn from $N(15,\sigma_{\text{outlier}}^{2})$, where $p_{\text{outlier}}\in\{0.1,0.2,0.3\}$ and $\sigma_{\text{outlier}}\in\{3,10\}$. In the $\sigma_{\text{outlier}}=3$ case, the generated outliers are likely to lie in the same range as the function values. Setting $\sigma_{\text{outlier}}=10$ constitutes a more difficult case where the outliers are unrelated to the function and are likely to lie outside of the function value range. We consider four scenarios with increasing outlier proportion and magnitude to study the performance of the considered noise models. These scenarios are summarized in table 2. The first scenario is the same one considered in Kuss, (2006) and serves as a baseline for the remaining scenarios which have an increasing proportion of extreme outliers. We fit the various GPR models to simulated datasets generated under various outlier scenarios and evaluate them on 10,000 noise-free test samples of eq. 25.
@@
Table 2: Summary of outlier scenarios considered in the second simulation study.
Boxplots of the RMSE, MAE, and NLPD for the different models and outlier scenarios are given in fig. 6. Note that the boxplots for GPR-CN and GPR-t tend to have a higher spread compared to GPR-G and GPR-L. This is likely because these methods have ELBOs that are not log-concave and are vulnerable to local optima. Furthermore, the NLPDs for GPR-t are approximated with Monte-Carlo samples. In all scenarios, the robust noise models (GPR-CN, GPR-t, GPR-L) outperform GPR-G in terms of all metrics. The RMSEs and MAEs are similar for the robust noise models (GPR-CN, GPR-t, GPR-L) in all scenarios. In the first scenario, the NLPDs are similar for GPR-CN and GPR-t but are slightly higher for GPR-L. This difference in NLPDs is exacerbated in the remaining scenarios, where the outliers are more extreme. This is likely because the Laplace distribution is not as heavy-tailed as the t-distribution and does not model outliers separately as in the CN distribution. As we increase the proportion of outliers, GPR-CN starts to outperform GPR-t in terms of NLPD. This suggests that GPR-CN is more adept at handling larger proportions of outliers than GPR-t. This may be due to the fact that outliers are identified and placed into its own separate component in GPR-CN so they don’t affect model fit for non-outliers. On the other hand, outliers are not explicitly isolated in GPR-t so they may affect the fit for non-outliers. As the proportion of outliers increases, we may expect the true latent function posterior to be multimodal, making the Gaussian variational distribution inadequate for approximating it. However, this does not seem to hinder predictive performance as can be seen from the results in the third and fourth scenario.
(a) $p_{\text{outlier}}=0.1$, $\sigma_{\text{outlier}}=3$
(b) $p_{\text{outlier}}=0.1$, $\sigma_{\text{outlier}}=10$
(c) $p_{\text{outlier}}=0.2$, $\sigma_{\text{outlier}}=10$
(d) $p_{\text{outlier}}=0.3$, $\sigma_{\text{outlier}}=10$
Figure 6: Boxplots of NLPD, RMSE, and MAE for second simulation study with varying $p_{\text{outlier}}$ and $\sigma_{\text{outlier}}$ across 200 simulated datasets.
@@ 5 Applications
In this section, we compare the numerical performances of the proposed algorithm with state-of-the-art literature on two data sets. One is the flight delay data, which has been adopted extensively in the Gaussian process literature, and the other is the ground magnetic perturbation data that we described in the Introduction. The flight delay data serves as a real-data evaluation of our model, and the ground magnetic perturbation data is of great scientific interest. We fit the four SVGPR models identified in section 4.2 on flight delays. For ground magnetic perturbations data ($\delta B_{H}$), we fit a dense artificial neural network and a GPR-CN model and compare their predictive performances with skill scores, RMSE, and focus in depth on the invterval coverage properties.
The model fitting procedure varies slightly between datasets. For the flights data, we train the four GPR models with 1000 inducing points for up to 30 iterations using the Adam algorithm implemented in PyTorch with an exponentially step size starting at 0.1 and a batch size of 256 (Kingma and Ba, 2014; Paszke et al., 2019). The inducing inputs were initialized to be a random sample of the training inputs. We used a validation set to monitor the validation NLPD and terminate training if it hasn’t decreased in 5 iterations. Input features are standardized by subtracting the training sample mean and dividing by the training sample standard deviation.
For $\delta B_{H}$ prediction, we use a more flexible hyperparameter setup, opting to allow the number of inducing points and batch size to vary and choosing optimal hyperparameters using the Bayesian Optimization package Nogueira, 14 . As the data is fully numeric and relatively small station by station, we use a large batch size on a GPU to speed up computation. The details of this can be found in 5.2.2.
@@ 5.1 Flight delays
The flight delays dataset consists of information about commercial flights in the US from January 2008 to April 2008 and was taken from an example in Hensman et al., (2013). It has become a standard benchmark dataset in the GPR literature for comparing scalable GPR methods due to its massive size and non-stationary nature. The original dataset curated by Hensman et al., (2013) contains 800,000 observations that were randomly sampled from a dataset containing around 2 million observations. To speed up computation, we reduced the size by randomly sampling 100,000 observations for training, and 20,000 each for validating and testing. The goal is to predict flight delay times (in minutes) using the eight features considered by Hensman et al., (2013): aircraft age, distance to travel, airtime, departure time, arrival time, day of the week, day of the month, and month. A histogram of flight delay times used for training is given in fig. 7. 99% of flights are only delayed for up to 166 minutes. However, there are a handful of flights that are delayed for more than 300 minutes. These delays are likely due to external factors such as weather that are not considered in our models. This motivates the use of robust models for analyzing this data.
@@ 5.2.1 Data Setup
Table 5: Summary statistics of GPR-CN and ANN models for the testing period of 2015-2016. Target coverage is 95%. Coverage is comparable in both storm and non-storm periods, but the IQR of interval length is generally either comparable or lower in the GPR-CN model.
All Test Data
@@
Storm Periods Only
@@
Following the convention made in the space weather community, we do not include lagged DeltaB values as inputs (Keesee et al., 2020; Upendran et al., 2022; Pinto et al., 2022). Therefore, we will ignore the temporal nature of this data and treat it as a conventional regression problem. Space physicists are typically most interested in the predictive performance during geomagnetic storms when $\delta B_{H}$ spikes, and so we report results across the entire prediction range as well as within storm performance.
@@ 5.2.2 Models
To fit the GPR-CN models, we use a Linux server with 32GB of RAM, a 16-core processor, and a 4GB NVidia GeForce 2060 GPU. As with the ANN, we use the Bayesian-Optimization library Nogueira, 14 to find optimal hyperparameters. While the GPU is not technically necessary, it should be noted that hyperparameter tuning on the CPU is a lengthy process, and GP models can be susceptible to the starting points of tuning parameters. The GPU enables running many more iterations of both model fitting and hyperparameter tuning, though it comes at a cost of model flexibility, namely by limiting the kernel setup. Due to memory constraints on the GPU, a periodic component could not be added to the existing kernel, and instead, seasonal features were added to the data to remedy this. Next, GPU training adds some complexity to the code primarily around data-type handling, though this is mitigated by the robust support of both the PyTorch and GPyTorch packages.
Though the GPU-training is optional, it vastly improved the number of iterations we were able to perform for both hyperparameter training and model testing, and is therefore recommended for any extensions to this research.
@@ 5.2.3 Results
In terms of accuracy, the ANN model performs slightly better than the GPR-CN model. Overall RMSE is lower on most stations, though both models are comparable. In addition to RMSE, we calculate Heidke and True Skill scores using station-specific deciles, as in Camporeale et al., (2020).
@@
5.2.4 Discussion
The GPR-CN model is a more conservative approach to modelling than the ANN. On average, the GPR-CN model under predicts when compared to the ANN. As a result, the conservatism of the GPR-CN model causes it to have diminished accuracy compared to the ANN across both the variance and bias of the errors (see table 9 in the appendix).
While this conservatism leads to lower accuracy, it yields better estimates of the variance for a given prediction. The ANN often overestimates the standard deviation of a given prediction to the point of creating arbitrarily large confidence intervals that achieve coverage by guessing an impossibly large range. This problem is especially pronounced in storm periods, but the GPR-CN method largely avoids this issue. Though the GPR-CN model yields worse accuracy, the largest errors are considerably smaller than the largest errors from the ANN. Across the majority of stations tested, the GPR-CN model has a much smaller 95th and 99th percentile of error, especially looking at high latitude stations (see table 6).
We have considered other uncertainty quantification methods such as MC dropout from Gal and Ghahramani, (2016). However, our data is a time series of extreme events, which introduces significant aleatoric (data) uncertainty due to high variability. MC Dropout primarily models epistemic (model) uncertainty and does not effectively capture aleatoric uncertainty. Even with an increased dropout rate of 0.7, the coverage rate with MC dropout remains poor (see table 8). By incorporating negative log-likelihood loss, our method models both the mean and variance, directly measuring aleatoric uncertainty.
Table 6: 95th and 99th percentile of errors for the ANN and GPR-CN models. GPR-CN consistently has lower errors at the top of the range for all stations but one.
Q95 Q99
Model ANN GPR-CN ANN GPR-CN
Station
ABK 6378.09 469.85 9948.65 1186.73
FRD 115.44 110.04 518.18 213.88
FRN 100.98 119.79 344.42 234.25
HRN 918.05 433.00 4288.03 713.55
IQA 1048.26 795.39 4367.02 1137.49
MEA 796.80 200.42 7301.83 564.96
NEW 121.76 99.35 480.85 193.53
OTT 112.96 96.77 446.79 175.52
PBQ 1794.57 525.56 8331.53 1141.94
WNG 110.40 90.96 390.30 165.83
YKC 1724.41 868.48 6991.68 1577.43
The primary advantage of the GPR-CN model is the interval length. Though the GPR-CN model sacrifices accuracy when compared to the ANN, it does so for the benefit of having better constrained prediction intervals. For nine stations of eleven, the median interval length was lower than the ANN, and in all stations, the GPR-CN model had lower maximum interval lengths for similar coverage (table 5, fig. 10). Further, as the RMSE of a particular station increases, the IQR of interval lengths increases as well, but this increase is less severe for the GPR-CN model (see fig. 8). This improvement to confidence intervals is especially pronouned examining the high RMSE (re: high latitude) stations during storm periods.
@@ 6 Conclusion
In this paper, we proposed a scalable inference algorithm for fitting sparse GPR models with contaminated normal noise. In section 4.1, we showed that the algorithm is able to correctly estimate the model hyperparameters when used to fit data simulated from the model itself. In section 4.2, we compared GPR models with different likelihoods and showed that our proposed model outperforms the others when there is a large proportion of outliers. In section 5, we applied our proposed algorithm to two real-world datasets and discussed strengths and weaknesses compared to GPR models with other robust likelihoods.
For the application area of geomagnetic perturbation forecasting, our improvements are two-fold. First, our model was able to provide more informative prediction intervals for the ground magnetic perturbations dataset when compared against other robust methods, including MC Dropout and an ANN with the target objective of estimating the parameters of a normal distribution. Next, our model consistently avoids large over-predictions, especially during periods of high geomagnetic activity. The implementation of a mixture-based likelihood for the Gaussian Process allows more stable estimates that are less influenced by outliers at comparable accuracy to state-of-the-art methods.